OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (5, 7) and (8, 22)?

OpenStudy (anonymous):

slope intercept form is y=mx+b so first find the slope by using the equation m=(y2-y1)/(x2-x1) and then plug in your known values into the equation y=mx+b to determine the value for b. Then leave your answer in terms of x and y. Your work should look like: m=(7-22)/(5-8)=5 y=5x+b -> 7=5(5)+b to find b=7/25 then plug in your known values: y=5x+(7/25)

OpenStudy (anonymous):

@cwrw238 help!!!!!